OpenStudy (anonymous):

can you help me with that : lim x2−−√3(x+1−−−−−√−x2−1−−−−−√) as x approch to infinity

OpenStudy (anonymous):

I'm confused as to what is happening in that equation.

OpenStudy (anonymous):

can you help me with that : lim\[\sqrt[3]{x^{2}}(\sqrt{x+1}-\sqrt{x^{2}-1})\] as x approaches to infinity @Bmannn13

OpenStudy (anonymous):

Sure!

OpenStudy (xapproachesinfinity):

let's look at the limit (sqrt{x+1}-sqrt{x^2-1}) since that one decides the whole thing

OpenStudy (xapproachesinfinity):

so what do you think we can do here?

OpenStudy (xapproachesinfinity):

any possible technique that we can use to get rid of indeterminate form

OpenStudy (anonymous):

I just got the answer, but someone else is helping already!

OpenStudy (xapproachesinfinity):

oh no answers :) just help

OpenStudy (xapproachesinfinity):

let the person do the work :)

OpenStudy (anonymous):

just write it please ! @Bmannn13

OpenStudy (xapproachesinfinity):

@M_lowreen it is not good for you ^_- you gotta do the work my dear

OpenStudy (xapproachesinfinity):

we can walk you through if you have the interest

OpenStudy (anonymous):

it's 00:20 am and i've studying for more than 6 hours.... can anyone help me please !

OpenStudy (anonymous):

That's true! I wasn't just going to give you the answer either.

OpenStudy (anonymous):

i know... i need just an idea !

OpenStudy (xapproachesinfinity):

does not matter! you still gotta do your own assignment by your own!

OpenStudy (anonymous):

Here, do this:

OpenStudy (anonymous):

\[d/dx(\sqrt{x+1}-\sqrt{x ^{2}-1})\]

OpenStudy (anonymous):

what does d/dx mean ? @Bmannn13

OpenStudy (anonymous):

Find the derivative.

OpenStudy (anonymous):

i haven't learned this yet! :/

OpenStudy (xapproachesinfinity):

okay so when we are faced with this sort of expression with radicals we can think of multiplying by "forget the word lol" just gonna give you an example \[(\sqrt{a}-\sqrt{b}) \frac{\sqrt{a}+\sqrt{b}}{\sqrt{a}+\sqrt{b}}\]

OpenStudy (xapproachesinfinity):

you can do that algebraic manipulation to get rid of indeterminate problem

OpenStudy (xapproachesinfinity):

surely you LEARNED THAT ONE ^_-

OpenStudy (xapproachesinfinity):

DARN MY MIND CAN'T RECALL THE NAME LOL

OpenStudy (anonymous):

ok, thanks ! I'll try it now :) :) @xapproachesinfinity

OpenStudy (xapproachesinfinity):

we do the same for our problem here

OpenStudy (xapproachesinfinity):

a is this case x+1 b is x^2-1

OpenStudy (xapproachesinfinity):

the game is to have some cancellations to make it more easier to study its limit

OpenStudy (anonymous):

I'm exchausted.... i;ve been studing since 5 pm..... tomorrow i'm going to take an advanced level exam @xapproachesinfinity

OpenStudy (xapproachesinfinity):

the word i was looking for it the conjugate so we multiply and divide by the conjugate darn the word didn;t want to come out hahaha

OpenStudy (xapproachesinfinity):

so what does that have to do with this! hahaha

OpenStudy (anonymous):

oww ... i know this !

OpenStudy (xapproachesinfinity):

I may sound hard on ya, but believe that won't help any good!

OpenStudy (xapproachesinfinity):

okay so we have \[\lim \sqrt{x+1}-\sqrt{x^2-1}=\lim (\sqrt{x+1}-\sqrt{x^2-1})\frac{\sqrt{x+1}+\sqrt{x^2-1}}{\sqrt{x+1}+\sqrt{x^2-1}}\]

OpenStudy (xapproachesinfinity):

first step

OpenStudy (xapproachesinfinity):

then \[=\lim\frac{(\sqrt{x+1}-\sqrt{x^2-1})(\sqrt{x+1}+\sqrt{x^2-1})}{\sqrt{x+1}+\sqrt{x^2-1}}\]

OpenStudy (xapproachesinfinity):

I just rewrite so it can appear to you as \[(a-b)(a+b)\]

OpenStudy (anonymous):

but \[\sqrt{x^{2}-1}=\sqrt{(x-1)(x+1)}\]

OpenStudy (xapproachesinfinity):

which you should now since you took algebra classes

OpenStudy (xapproachesinfinity):

nope

OpenStudy (xapproachesinfinity):

\[=\lim \frac{(\sqrt{x+1})^2-(\sqrt{x^2-1})^2}{\sqrt{x+1}+\sqrt{x^2-1}}\]

OpenStudy (anonymous):

yeah,, i know right... up to here ... but then?

OpenStudy (xapproachesinfinity):

you should already know that \[(a-b)(a+b)=a^2-b^2\] i used this since you are in calc class this has to be stuck in your memory

OpenStudy (anonymous):

i know ittt

OpenStudy (xapproachesinfinity):

earlier you said you can't do it now you saying you know???

OpenStudy (xapproachesinfinity):

then why you didn't do it?

OpenStudy (anonymous):

i said i know this algebra formulas ,,, but they confuse me all

OpenStudy (xapproachesinfinity):

if they confuse then you don't know them yet lol

OpenStudy (xapproachesinfinity):

just kidding :)

OpenStudy (anonymous):

Problem solved *___* lol

OpenStudy (xapproachesinfinity):

so back to your problem \[=\lim \frac{x+1-x^2+1}{\sqrt{x+1}+\sqrt{x^2-1}}\]

OpenStudy (xapproachesinfinity):

simplify more

OpenStudy (anonymous):

what if i factorise at the denominator \[\sqrt{x+1}(1+\sqrt{x-1})\] ? @xapproachesinfinity

OpenStudy (xapproachesinfinity):

how is that possible ? cannot be done what you did has no meaning try to distribute and see if you get back to that denominator

OpenStudy (xapproachesinfinity):

i was talking about the top simplify more i meant just add 1+1 to get 2+x-x^2

OpenStudy (xapproachesinfinity):

at this level it is still indeterminate form oo/oo which is no good

OpenStudy (xapproachesinfinity):

so we have to do something more to bring it to a more simplified form

OpenStudy (anonymous):

i know ,, sooo saaadddd

OpenStudy (xapproachesinfinity):

hehe \[\lim \frac{2+x-x^2}{\sqrt{x+1}+\sqrt{x^2-1}}=\lim\frac{2/x+1-x}{(\sqrt{1/x+1/x^2}+\sqrt{1-1/x^2})}\]

OpenStudy (xapproachesinfinity):

i factored out x top and bottom

OpenStudy (xapproachesinfinity):

now if you tried the evaluate the limit is should be easy

OpenStudy (xapproachesinfinity):

2/x goes to zero 1 to 1 -x goes to -oo so the top by it self goes to -oo ======== the bottom goes to 1 so we have -oo/1 which is -oo

OpenStudy (anonymous):

***Hugs*** :)))))) @xapproachesinfinity

OpenStudy (anonymous):

you made my day :D

OpenStudy (xapproachesinfinity):

all that work can be done in one scoop if you know about l'hopitals rule

OpenStudy (xapproachesinfinity):

no problem

OpenStudy (anonymous):

you may be right.. but it's the first time we work with limits :/ @xapproachesinfinity